OpenStudy (anonymous):

solve the system of equations : what is the x - value for the solution ? 2x + 7y =3 and x = 1 - 4y A. -1 B. 5 C. 4 D. 3 ?

OpenStudy (silverlark2):

2x+7y -(1) x= 1-4y -(2) Sub (2) into (1): 2(1-4y) + 7y = 3 2-8y+7y = 3 -y= 1 y= -1 Hence sub y=-1 into (2): x= 1-4(-1) x=5 Ans: B